Noluvpablo:

Solve for x: 5(x + 3) = 4(x − 3) −18 −27 −3 −6

Noluvpablo:

need help

martai:

ok i got u

martai:

give me a mim

martai:

what do u think it is ?

404:

Solve for x by simplifying both sides of the equation, then isolating the variable.

YRJ8498:

Distribute 5(𝑥+3)=4(𝑥−3) 5𝑥+15=4(𝑥−3)

YRJ8498:

after that you 5𝑥+15=4(𝑥−3) 5𝑥+15=4𝑥−12

YRJ8498:

so ditribute

YRJ8498:

then Subtract 15 from both sides of the equation 5𝑥+15=4𝑥−12 5x+15=4x-12−15

YRJ8498:

after that you should get the solution which is 𝑥=-27

YRJ8498:

Did this help?